OpenStudy (anonymous):

divide a polynomial by a binomial. How do you know if the binomial is a factor of the polynomial? Create a sample problem that has a binomial which IS a factor of the polynomial being divided, and another problem that has a binomial which is NOT a factor of the polynomial being divided. PLS help i'm so lost..

OpenStudy (anonymous):

HEY

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

BYE

OpenStudy (anonymous):

ok bye o.O

OpenStudy (campbell_st):

if you have a binomial (x + a) then if you substitute x = -a into the polynomial P(x) then if P(-a) = 0, (x + a) is a factor this is called the factor theorem. as a simple example is (x - 2) a factor of \[P(x) = x^2 +5x -14\] then evaluate P(2) \[P(2) = (2)^2 + 5(2) - 14\] P(2) = 0 so a factor

OpenStudy (anonymous):

im still lost and is that considered an example problem

OpenStudy (campbell_st):

yes it it... the basics are if you have a binomial.... (x + a): substitute x = -a (x -a): substitute x = a into the polynomial.... if the answer is zero.. then you have a factor.